# Conformal prediction for Deep Classifier via Label Ranking

### v1

@@ 6 Related work
Conformal prediction is a statistical framework characterized by a finite-sample coverage guarantee. It has been utilized in various tasks including regression (Lei & Wasserman, 2014; Romano et al., 2019), classification (Sadinle et al., 2019), structured prediction (Bates et al., 2021), Large-Language Model (Kumar et al., 2023; Ren et al., 2023) and so on.
@@
The primary focal points of CP are reducing prediction set size and enhancing coverage rate. Strategies to reduce the set size can be roughly split into the following two branches. The first approach involves leveraging post-hoc technologies (Romano et al., 2020; Angelopoulos et al., 2021a; Ghosh et al., 2023). There exist others concentrate on unique settings such as federated learning (Lu et al., 2023) and multi-label problem (Cauchois et al., 2020; Fisch et al., 2022; Papadopoulos, 2014), outlier detection (Bates et al., 2023; Chen et al., 2023; Guan & Tibshirani, 2022). Most existing post-hoc methods construct the non-conformity score based on unreliable probability values, leading to suboptimal performance. Different from previous post-hoc methods, we show that probability value is not necessary in non-conformity scores and design an effective method to remove the probability value while retaining uncertainty information.
Another avenue of research focuses on developing new training algorithms to reduce the average prediction set size (Colombo & Vovk, 2020; Chen et al., 2021; Stutz et al., 2022; Einbinder et al., 2022b; Bai et al., 2022; Fisch et al., 2021). Those training methods are usually computationally expensive due to the model retraining. Additionally, there is a growing number of work dedicated to enhancing coverage rate (Vovk, 2012; Shi et al., 2013; Löfström et al., 2015; Ding et al., 2023), including efforts to maintain the marginal coverage rate by modifying the assumption of exchangeability to accommodate factors such as adversaries (Gendler et al., 2021), covariate shifts (Tibshirani et al., 2019), label shifts (Podkopaev & Ramdas, 2021) and noisy labels (Einbinder et al., 2022a; Sesia et al., 2023). In this study, SAPS not only lessens the prediction sets, but also broadly enhances the conditional coverage rate and adaptation of prediction sets.
@@
In this paper, we present SAPS, a simple alternative CP algorithm that generates smaller prediction sets. By integrating the label rank, SAPS effectively mitigates the negative effect of small probability, resulting in a stable prediction set. The extensive experiments show that SAPS can improve the conditional coverage rate and adaptation while maintaining a small prediction set. This method can be easily applied to any pre-trained classifiers. We hope that our insights inspire future research to leverage label ranking information for conformal prediction.
Appendix A Proof of Theorem 1
@@
The non-conformity scores of examples for which the rank of ground-truth labels is $r$ can be computed by
$S(\bm{x},y,u;\hat{\pi})=r-u.$
By the above equation, we can observe that the scores of these examples are distributed uniformly on $[(r-1),r]$. For simplicity, let $a_{k}=A_{r}-A_{r-1}$ for $r>1$ and $a_{1}=A_{1}$. Moreover, suppose that there are infinite examples in the calibration set. We assume that $k$ satisfies $A_{k}\geq 1-\alpha>A_{k-1}$. Given that $\sum_{i=1}^{k-1}a_{i}=A_{k-1}$, the $1-\alpha$ quantile of scores for calibration examples, i.e., the calibrated threshold $\tau$, can be obtained by
$*.\tau=(k-1)+\frac{1-\alpha-\sum^{k-1}_{i=1}a_{i}}{a_{k}}*.$
Based on the definition of prediction set in Eq. 3, the prediction set of test example $(\bm{x},y)$ is equal to
$\mathcal{C}(\bm{x})=\{y|S(\bm{x},y,u;\hat{\pi})\leq\tau,y\in\mathcal{Y}\}.$ (9)
When $y$ satisfies the inequation $o(y,\hat{\pi}(\bm{x}))<k$, $S(\bm{x},y,u;\hat{\pi})$ must be smaller than $\tau$. Thus, the least size of sets is $k-1$. But for $y$ satisfying $o(y,\hat{\pi}(\bm{x}))=k$, we can have
$S(\bm{x},y,u;\hat{\pi})$ $\leq\tau$
$k-u$ $*\displaystyle\leq(k-1)+\frac{1-\alpha-\sum^{k-1}_{i=1}a_{i}}{a_{k}}*$
$1-\frac{1-\alpha-\sum^{k-1}_{i=1}a_{i}}{a_{k}}$ $\leq u$
Thus, there exists a probability of $\frac{1-\alpha-\sum^{k-1}_{i=1}a_{i}}{a_{k}}$ such that $S(\bm{x},y,u;\hat{\pi})\leq\tau$. Finally, the expected value of the set size is
$\mathbb{E}_{u\sim U}[|\mathcal{C}(\tilde{\bm{x}})|]=k-1+\frac{1-\alpha-\sum^{k-1}_{i=1}a_{i}}{a_{k}}=(k-1)+\frac{1-\alpha-A_{k-1}}{A_{k}-A_{k-1}}$ (10)
Appendix B Proof of Proposition 1
We first introduce a coverage guarantee theorem as shown in the following:
Theorem 2.
(Conformal calibration coverage guarantee,(Angelopoulos et al., 2021b)). Suppose $\left(\bm{x}_{i},y_{i},u_{i}\right)_{i=1,\ldots,n}$ and $\left(\bm{x}_{n+1},y_{n+1},u_{n+1}\right)$ are i.i.d. and let $\mathcal{C}_{1-\alpha}(\bm{x},u)$ be a set-valued mapping satisfying the nesting property in Eq. 2. Then for a confidence level $\alpha$, we have the following coverage guarantee:
$P\left(y_{n+1}\in\mathcal{C}_{1-\alpha}\left(\bm{x}_{n+1},u_{n+1}\right)\right)\geq 1-\alpha$
Then, we prove the finite-sample coverage guarantee of SAPS.
@@ Appendix B Proof of Proposition 1
Let $\alpha_{1}>\alpha_{2}$ and $S_{i}$ represents the score of example $(\bm{x}_{i},y_{i}),\forall i\in\{1,\dots,n\}$. Then, the calibrated threshold $\tau_{1}$ for $\alpha_{1}$ is $1-\alpha_{1}$ quantile of $\{S_{i}\}_{i=1}^{n}\cup\{\infty\}$. It is easy to have $\tau_{1}\leq\tau_{2}$. So, for a label $y\in\mathcal{Y}$, if $S(\bm{x}_{n+1},y,u;\hat{\pi})<\tau_{1}$, $S(\bm{x}_{n+1},y,u;\hat{\pi})<\tau_{2}$. Thus, $\mathcal{C}_{1-\alpha_{1}}\left(\bm{x}_{n+1},u_{n+1}\right)\subseteq\mathcal{C}_{1-\alpha_{1}}\left(\bm{x}_{n+1},u_{n+1}\right)$. After given the nesting property of $\mathcal{C}_{1-\alpha}\left(\bm{x}_{n+1},u_{n+1}\right)$, based on the Theorem 2, the following coverage guarantee can be obtained:
$P\left(y_{n+1}\in\mathcal{C}\left(\bm{x}_{n+1},u_{n+1}\right)\right)\geq 1-\alpha.$
Appendix C Proof of Proposition 2
We first introduce Theorem 3 describing the distribution of non-conformity scores in APS, then prove Lemma 1 for bounds of softmax probability. Finally, we give the proof of Proposition 2.
Theorem 3.
(Einbinder et al., 2022b) The distribution of the non-conformity scores $S$ in equation 4 is uniform conditional on $\bm{x}$ if $\hat{\pi}=\pi$. That is, $\mathbb{P}[\tilde{S}(\bm{x},y,u;\pi)\leq\beta\mid X=\bm{x}]=\beta$ for all $\beta\in(0,1)$, where $(\bm{x},y)$ is a random sample from $\mathcal{P}$, and $u\sim$ Uniform$[0,1]$ independent of everything else.
Lemma 1.
Given a discrete probability $\{p_{i}\}_{i=1}^{K}$ and $p_{i}>0,\forall i\in\{1,\dots,K\}$. The ordered probability list is denoted by $p_{(1)}>=p_{(2)},\dots>=p_{(K)}$. Then, we have that
$\frac{1}{K}<=p_{(1)}\leq 1,$
and
$0<p_{(k)}\leq\frac{1}{k},k\in\{2,\dots,K\}.$
Proof.
We have known that $\sum_{i=1}^{K}p_{i}=1$. Supposed that $p_{(1)}<\frac{1}{K}$. We can have that for $k\geq 2$,
$p_{(k)}\leq p_{(1)}<\frac{1}{K}.$
Thus,
$\sum_{i=1}^{K}p_{i}<1.$
Finally, there exists a contradiction. It is obviously that $p_{(1)}\leq 1$. We can conclude that $\frac{1}{K}<=p_{(1)}\leq 1$.
Supposed that $p_{(1)}>\frac{1}{k}$. We can have that for $\tilde{k}<k$,
$\sum_{i=1}^{k}p_{(i)}>k\cdot\frac{1}{k}=1.$
The above equation is contrastive to $\sum_{i=1}^{K}p_{i}=1$. It is obvious that $0<p_{(k)}$. Finally, we can conclude that $p_{(k)}\leq\frac{1}{k}$. ∎
Next, we prove Proposition 2.
Proof.
Let $S_{i}(\cdot)$ and $\tilde{S}_{i}(\cdot)$ denote the non-conformity scores of SAPS and APS, respectively. Since $\hat{\pi}$ is well-calibrated, i.e., $\hat{\pi}=\pi$, for any point $(\bm{x}_{i},y_{i})$ from calibration set, $S_{i}(\cdot)$ and $\tilde{S}_{i}(\cdot)$ are reduced to $u\cdot~{}\pi_{(1)}(\bm{x}_{i})$. Then,
$\mathbb{E}_{u\sim[0,1]}[S(x_{i},y_{i},u)]=\mathbb{E}_{u\sim[0,1]}[\tilde{S}(\bm{x}_{i},y_{i},u)]$
Based on Theorem 3, $s_{i}$ and $\tilde{s}_{i}$ are uniform condition $\bm{x}$, and we can get that $\tau=\tilde{\tau}=1-\alpha$. Next, considering two scenarios, i.e., $\pi_{(1)}(\bm{x})\geq\tau$ and $\pi_{(1)}(\bm{x})<\tau$.
If $\pi_{(1)}(\bm{x})\geq\tau$, the probability of $u\cdot\pi_{(1)}(\bm{x}_{i})\leq\tau$ is $\frac{\tau}{\pi_{(1)}(\bm{x})}$. Thus,
$\mathbb{E}_{u\sim[0,1]}\{|\tilde{\mathcal{C}}(\bm{x},u)|\}=\mathbb{E}_{u\sim[0,1]}\{|\mathcal{C}(\bm{x},u)|\}=\frac{\tau}{\pi_{(1)}(\bm{x})}$
If $\pi_{(1)}(\bm{x})<\tau$, without loss of generality, supposed $\sum_{k=1}^{m}\pi{(k)}<\tau$ and $\sum_{k=1}^{m+1}\pi_{(k)}>\tau$.
$\mathbb{E}_{u\sim[0,1]}\{|\tilde{\mathcal{C}}(\bm{x},u)|\}=m+\frac{\tau-\sum_{k=1}^{m}\pi{(k)}}{\pi_{(m+1)}(\bm{x})}$
But for SAPS, the analysis is more complex. Since $\pi_{(1)}(\bm{x})>\frac{1}{K}$, $\pi_{(1)}(\bm{x})+\lambda>1$
$\mathbb{E}_{u\sim[0,1]}\{|\mathcal{C}(\bm{x},u)|\}=1+\frac{\tau-\pi_{(1)}(\bm{x})}{\lambda}.$
When $m\geq 2$, with $\frac{1}{\lambda}\leq 1-\frac{1}{K}$ and $\pi_{(1)}(\bm{x})\in[\frac{1}{K},1]$ proposed in Lemma 1,
$\mathbb{E}_{u\sim[0,1]}\{|\mathcal{C}(\bm{x},u)|\}=1+\frac{\tau-\pi_{(1)}(\bm{x})}{\lambda}\leq 1+\frac{\tau-\frac{1}{K}}{1-\frac{1}{K}}<2<\mathbb{E}_{u\sim[0,1]}\{|\tilde{\mathcal{C}}(\bm{x},u)|\}.$
When $m=1$, with $\lambda>\pi_{(2)}(\bm{x})$ in Lemma 1, we can get
$1+\frac{\tau-\pi_{(1)}(\bm{x})}{\lambda}<1+\frac{\tau-\pi_{(1)}(\bm{x})}{\pi_{(2)}(\bm{x})}$
Finally, $\mathbb{E}_{u\sim[0,1]}\{|\mathcal{C}(\bm{x},u)|\}\leq\mathbb{E}_{u\sim[0,1]}\{|\tilde{\mathcal{C}}(\bm{x},u)|\}$. ∎
Appendix D Detailed results for coverage rate and set size
@@
In this section, we report the detailed results of coverage rate and set size on different datasets in Table 3,4, and 5. The median-of-means for each result is reported over ten different trials. The average results of multiple models have been reported in the main paper.
@@
Appendix E ESCV on CIAFR-10 and CIFAR-100
In this section, we report the results of ESCV on CIFAR-10 and CIFAR-100 with $\alpha=0.1$. From Figure 5 and 5, we can observe that the SAPS still can get smaller ESCV than RAPS. Moreover, as the scale of the classification task increases, the ESCV for RAPS also increases significantly, while SAPS maintains a low value. For example, on CLIP, the ESCV for RAPS on ImageNet is higher than that for CIFAR-100. But, for both ImageNet and CIFAR-100, the ESCV of SAPS remains around $0.35$.
[…]
Figure 5: ESCV with different models on CIFAR-10 and CIFAR-100 with $\alpha=0.1$.
Appendix F adaptation
In this section, we report detailed results of adaptation for four models with different architectures on ImageNet. In Table 6, for all models, SAPS achieves a smaller set size for easy examples than SAPS, but a larger set size for hard examples.
Table 6: Size conditional on difficulty. We report the size of RAPS and SAPS with ResNet152, VGG16, Inception, and ViT on ImageNet. The term ”Difficulty” refers to the relative ranking of the probability estimates of the true label class. The ”Count” corresponds to the number of examples falling within each difficulty category.
ResNet152 VGG16 Inception ViT
Count Size Count Size Count Size Count Size
Difficulty RAPS SAPS RAPS SAPS RAPS SAPS RAPS SAPS
1 15696 1.87 1.57 14316 3.26 2.57 13905 4.72 3.32 16207 1.56 1.36
2-3 2560 2.79 2.85 3019 4.63 4.82 3058 6.63 6.81 2404 2.55 2.52
4-6 712 3.06 3.29 1013 5.11 5.49 997 7.36 8.19 598 2.78 2.78
7-10 345 3.10 3.39 455 5.23 5.72 584 7.47 8.57 267 2.81 2.96
11-100 615 3.19 3.51 1057 5.34 6.01 1116 7.77 9.09 452 3.05 3.26
101-1000 72 3.50 3.83 140 5.50 6.12 340 7.51 8.85 72 2.76 2.72
Appendix G SAPS vs. RAPS($k_{r}=1$)
To further understand the influence of excluding the probability value, we conduct an experiment for RAPS($k_{r}=1$), which is similar to SAPS on ImageNet, CIFAR-100, and CIFAR-10 with $\alpha=0.1$. The median results on ten trials are reported in Table 7, 8 and 9.

### v2

@@ 6 Related work
Conformal prediction (CP) is a statistical framework characterized by a finite-sample coverage guarantee (Vovk et al., 2005). It has been utilized in various tasks including regression  (Lei & Wasserman, 2014; Romano et al., 2019), classification  (Sadinle et al., 2019), structured prediction  (Bates et al., 2021), Large-Language Model  (Kumar et al., 2023; Ren et al., 2023; Quach et al., 2023; Su et al., 2024), Diffusion Model (Teneggi et al., 2023; Horwitz & Hoshen, 2022), robots control (Wang et al., 2023), graph neural network (Wijegunawardana et al., 2020; Clarkson, 2023; Song et al., 2024) and so on. Conformal prediction also is deployed in other applications, such as human-in-the-loop decision making (Straitouri et al., 2023; Cresswell et al., 2024), automated vehicles (Bang et al., 2024), and scientific machine learning (Moya et al., 2024; Podina et al., 2024). Besides the inductive conformal prediction (Vovk et al., 2005), most CP methods follow an inductive conformal framework (or split conformal prediction) with a held-out calibration set to cope with the relative computational inefficiency of conformal predictors (Papadopoulos et al., 2002a, b; Vovk, 2002), as in the main paper (Lei et al., 2015). Additionally, there are various variants of conformal predictors based on cross-validation (Vovk, 2015) or jackknife (i.e., leave-one-out) (Barber et al., 2021). The primary focal points of CP are reducing prediction set size and enhancing coverage rate.
Inefficiency.
@@
Strategies to improve the efficiency of prediction sets can be roughly split into the following two branches. The first avenue of research focuses on developing new training algorithms to reduce the average prediction set size (Bellotti, 2021; Colombo & Vovk, 2020; Chen et al., 2021; Stutz et al., 2022; Einbinder et al., 2022; Bai et al., 2022; Fisch et al., 2021; Yang & Kuchibhotla, 2021). Those training methods are usually computationally expensive due to the model retraining and the high complexity of optimization. Thus, CP has also been studied in the context of Early Stopping (Liang et al., 2023), allowing model selection leading to compact prediction sets while keeping coverage guarantees. The other avenue involves leveraging post-hoc technologies, such as novel score functions (Romano et al., 2020; Angelopoulos et al., 2021; Ghosh et al., 2023) and post-hoc learning method (Xi et al., 2024). There exist others concentrating on unique settings such as federated learning (Lu et al., 2023; Plassier et al., 2023), multi-label problem (Cauchois et al., 2021; Fisch et al., 2022; Papadopoulos, 2014), outlier detection (Bates et al., 2023; Guan & Tibshirani, 2022), and out-of-distribution detection (Chen et al., 2023; Novello et al., 2024). Most existing post-hoc methods calculate the non-conformity score based on unreliable probability values, leading to sub-optimal performance. In contrast to previous post-hoc methods, we show that probability value is not necessary for non-conformity scores and design an effective method to remove the probability value while retaining uncertainty information.
Validity of coverage rate. There is an increasing amount of research focused on improving coverage rates(Shi et al., 2013; Löfström et al., 2015; Ding et al., 2023), including efforts to maintain the marginal coverage rate by modifying the assumption of exchangeability to accommodate factors such as adversaries (Gendler et al., 2021; Kang et al., 2024), covariate shifts (Tibshirani et al., 2019; Deng et al., 2023; Xu et al., 2024), label shifts (Podkopaev & Ramdas, 2021; Plassier et al., 2023) and noisy labels (Feldman et al., 2023; Sesia et al., 2023). Moreover, conformal prediction algorithms pursue conditional coverage guarantee (Vovk, 2012), including training-conditional validity and object-conditional validity. Specifically, some methods acquire the training-conditional coverage guarantee which would ensure that most draws of the training data set to result in valid marginal coverage on future test points (Bian & Barber, 2023; Pournaderi & Xiang, 2024). There are also group-conditional conformal prediction methods that guarantee performance on all groups of individuals in the population (Javanmard et al., 2022; Gibbs et al., 2023; Melki et al., 2023). Although it is impossible to obtain conditional coverage (Foygel Barber et al., 2021), THR (Sadinle et al., 2019) shows that efficient class-conditional coverage is possible. For example, Clustered CP (Ding et al., 2023) utilizes the taxonomy of the label space to improve the class-conditional coverage rate and $k$-Class-conditional CP calibrates the class-specific scores threshold depending on the top-$k$ error. As discussed in Sec.4.2, SAPS broadly enhances the conditional coverage rate of prediction sets while maintaining small prediction sets.
@@ 7 Conclusion
In this paper, we present SAPS, a simple and effective conformal prediction score function that discards almost all the probability values except for the maximum softmax probability. By highlighting the information on label ranking, SAPS effectively mitigates the negative effect of tail probabilities, resulting in small and stable prediction sets. Additionally, we provide a key insight that the value of softmax probability might be redundant information for conformal prediction. Extensive experiments show that SAPS can improve the conditional coverage rate while maintaining a small prediction set. Our method is straightforward to implement with any classifier and can be easily adopted in various practical settings. We expect that out method can inspire future research to emphasize the label ranking information for conformal prediction.
8 Limitations
The limitation of our method is mainly on the hyperparameter representing the weight of label ranking that requires to be tuned with a validation dataset (same as RAPS). About future direction, we expect to enhance the conditional coverage of these conformal prediction algorithms on large-scale datasets, where all CP methods achieve worse conditional coverage rates than those of small-scale datasets.
Acknowledgements
This research is supported by the Shenzhen Fundamental Research Program (Grant No. JCYJ20230807091809020). Yue Qiu is supported by the National Natural Science Foundation of China (Grant No. 12101407), the Chongqing Entrepreneurship and Innovation Program for Returned Overseas Scholars (Grant No. CX2023068), and the Fundamental Research Funds for the Central Universities (Grant No. 2022CDJJCLK-002, 2023CDJXY-042). We gratefully acknowledge the support of the Center for Computational Science and Engineering at the Southern University of Science and Technology, as well as the Key Laboratory of Nonlinear Analysis and its Applications at Chongqing University, under the Ministry of Education, China, for our research.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Proof of Proposition 3.2
@@ Acknowledgements
Let $\hat{\pi}$ be a classifier. $(\boldsymbol{x}_{1},y_{1})$ and $(\boldsymbol{x}_{2},y_{2})$ represent two instances sampled from the joint data distribution $\mathcal{P}_{\mathcal{XY}}$, and $u_{1},u_{2}$ are sampled from $U[0,1]$. It is easy to get that for any different positive numbers $\gamma_{1}$ and $\gamma_{2}$, $\gamma_{1}\cdot[o(y_{1},\hat{\pi}(\boldsymbol{x}_{1}))-1+u_{1}]\leq\gamma_{1}\cdot[o(y_{2},\hat{\pi}(\boldsymbol{x}_{2}))-1+u_{2}]$ if and only if $\gamma_{2}\cdot[o(y_{1},\hat{\pi}(\boldsymbol{x}_{1}))-1+u_{1}]\leq\gamma_{2}\cdot[o(y_{2},\hat{\pi}(\boldsymbol{x}_{2}))-1+u_{2}]$. According to Lemma 3.1, we can get that $S_{cons}(\boldsymbol{x},y;\hat{\pi},\gamma_{1})$ and $S_{cons}(\boldsymbol{x},y;\hat{\pi},\gamma_{2})$ could provide the same prediction sets. ∎
Appendix B Proof of Theorem 3.3
@@
For simplicity, let $a_{r}=A_{r}-A_{r-1}$ for $r\geq 1$ and $A_{0}=0$. $a_{r}$ represents the rate of examples in the calibration set whose ground-truth labels’ rank is $r$. The non-conformity scores of examples whose rank of ground-truth labels are $r$ can be computed by
$S_{cons}(\boldsymbol{x},y,u;\hat{\pi})=r-u,$
where $u\sim U[0,1]$. Thus, we can observe that the scores of these examples are uniformly distributed on $[r-1,r]$. Scores of examples whose rank of ground-truth labels are less than or equal to $r$ are no more than $r$. In other words, the rate of examples in the calibration set with scores lower than $r$ is $A_{r}$. Given that $k$ satisfies $A_{k}\geq 1-\alpha>A_{k-1}$, the $1-\alpha$ quantile of scores for the calibration set, i.e., the calibrated threshold $\tau$, is located in the interval $[k-1,k]$. Since the scores in the interval $[k-1,k]$ are uniformly distributed and the rate of examples whose scores in $[k-1,k]$ is $a_{k}$, $\tau$ is equivalent to the $\frac{1-\alpha-A_{k-1}}{a_{k}}$ quantile of $[k-1,k]$. Formally, we can get the value of $\tau$ by
$\tau=(k-1)+\frac{1-\alpha-A_{k-1}}{a_{k}}.$
Based on the definition of prediction set in Equation 3, the prediction set of a test instance $\boldsymbol{x}$ is equal to
$\mathcal{C}_{1-\alpha}(\boldsymbol{x},u)=\{y\in\mathcal{Y}:o(y,\hat{\pi}(\boldsymbol{x}))-u\leq\tau\}.$ (11)
Next, we analyze the prediction set $\mathcal{C}_{1-\alpha}(\boldsymbol{x},u)$. When $y\in\mathcal{Y}$ satisfies the inequation $o(y,\hat{\pi}(\boldsymbol{x}))\leq k-1$, $S_{cons}(\boldsymbol{x},y,u;\hat{\pi})$ must be smaller than $\tau$. Thus, the size of sets is at least $k-1$. But for $y$ satisfying $o(y,\hat{\pi}(\boldsymbol{x}))=k$, we can have
$o(y,\hat{\pi}(\boldsymbol{x}))-u$ $\leq\tau$
$k-u$ $\leq(k-1)+\frac{1-\alpha-A_{k-1}}{a_{k}}$
$1-\frac{1-\alpha-A_{k-1}}{a_{k}}$ $\leq u$
Thus, there exists a probability of $\frac{1-\alpha-A_{k-1}}{a_{k}}$ such that $S(\boldsymbol{x},y,u;\hat{\pi})\leq\tau$. Finally, the expected value of the set size for the test example $\boldsymbol{x}$ is
$\mathbb{E}_{u\sim U}[|\mathcal{C}_{1-\alpha}(\boldsymbol{x},u)|]=k-1+\frac{1-\alpha-A_{k-1}}{a_{k}}=k-1+\frac{1-\alpha-A_{k-1}}{A_{k}-A_{k-1}}$ (12)
Appendix C Detailed results for coverage rate and set size
@@
In this section, we report the detailed results of coverage rate and set size on different datasets in Table 5,6, and 7. The median-of-means for each result is reported over ten different trials. The average results of multiple models have been reported in the main paper.
@@
Appendix D ESCV on CIAFR-10 and CIFAR-100
In this section, we report the results of ESCV on CIFAR-10 and CIFAR-100 with $\alpha=0.1$. From Figure 4(a) and 4(b), we can observe that the SAPS still can get smaller ESCV than RAPS. Moreover, as the scale of the classification task increases, the ESCV for RAPS also exhibits a marked increase, whereas SAPS consistently demonstrates a low value. For example, with the CLIP model, the ESCV for RAPS on ImageNet exceeds that observed for CIFAR-100. But, for both ImageNet and CIFAR-100, the ESCV of SAPS remains around $0.35$.
((a)) CIFAR-10
((b)) CIFAR-100
Figure 4: ESCV for different models on CIFAR-10 and CIFAR-100 with $\alpha=0.1$.
Appendix E SAPS vs. RAPS($k_{r}=1$)
To further understand the influence of excluding the probability value, we conduct an experiment for RAPS($k_{r}=1$), which is similar to SAPS on ImageNet, CIFAR-100, and CIFAR-10 with $\alpha=0.1$. The median results on ten trials are reported in Table 8, 9 and 10.